OpenStudy (anonymous):

Which of the following gives a recursive representation with n > 2 of the partial sum of the geometric series below? 4 + 8 + 16 + 32..... A. Sn = 4 + 2Sn-1 B. Sn = 4 + (1/2)Sn-1 C. Sn = 4 + Sn-1 D. Sn = 2Sn-1

OpenStudy (anonymous):

@Annie7077 help?

OpenStudy (anonymous):

what is Sn

OpenStudy (anonymous):

S small n

OpenStudy (anonymous):

whatever that means

jimthompson5910 (jim_thompson5910):

It's the nth partial sum

jimthompson5910 (jim_thompson5910):

I think the best way to do this is to just check each option

jimthompson5910 (jim_thompson5910):

For instance, option D S1 = 4 S2 = 4+8 = 12 notice how S2 is NOT equal to 2*S1, so option D is out

OpenStudy (anonymous):

ok where is n in this equatio?

jimthompson5910 (jim_thompson5910):

n refers to the index number of each term (n = 1 is the first term, n = 2 is the second, etc)

OpenStudy (anonymous):

yes but where is it in this equation? where to substitute it in?

jimthompson5910 (jim_thompson5910):

S1 is found by adding the first term. Basically S1 is the first term itself

jimthompson5910 (jim_thompson5910):

S2 is found by adding the first two terms S2 = 4+8 = 12

jimthompson5910 (jim_thompson5910):

S3 is found by adding the first three terms S3 = 4+8+16 = 28

jimthompson5910 (jim_thompson5910):

etc etc

jimthompson5910 (jim_thompson5910):

hopefully you see why D is false?

OpenStudy (anonymous):

@jim_thompson5910 so a?

jimthompson5910 (jim_thompson5910):

let's find out

jimthompson5910 (jim_thompson5910):

S2 = 4 + 2S1 12 = 4 + 2*4 12 = 12 ... works ------------------------------------------------------- S3 = 4 + 2*S2 28 = 4 + 2*12 28 = 28 ... works ------------------------------------------------------- It turns out the others work as well, so you are correct

jimthompson5910 (jim_thompson5910):

I put that last statement in because there is no upper limit on what to check, so you'd be at it forever.

OpenStudy (anonymous):

@jim_thompson5910 thank you!

jimthompson5910 (jim_thompson5910):

you're welcome